Use the comparison method to solve the following system of equations:
(1) 2x - 3y = 5. (2) 2x + 5y = -3

Answer by ikleyn(52788)

.
Use the comparison method to solve the following system of equations:
(1) 2x - 3y = 5. (2) 2x + 5y = -3
------------------------------------------

Distract the second equation from the first one.

In this way you exclude x and get the unique equation for y.